OpenStudy (anonymous):

Factor x^3-11x^2-4x+224 completely long division if (x-8) is a factor

OpenStudy (anonymous):

x^3-11x^2-4x+224 = (x-8)(g(x)) g(x) = (x^3-11x^2-4x+224)/(x-8) . . . . x^2-3x-28 = g(x) . . . .------------------- x-8 | x^3-11x^2-4x+224 . . . . . x^3-8x^2 . . . . . --------- . . . . . . . . .-3x^2-4x . . . . . . . . .-3x^2+24x . . . . . . . . .---------- . . . . . . . . . . . . . .-28x+224 . . . . . . . . . . . . . .-28x+224 . . . . . . . . . . . . . .--------- . . . . . . . . . . . . . . . . . . . .0 x^3-11x^2-4x+224 = (x-8)(x^2-3x-28) = (x-8)(x-7)(x+4)

OpenStudy (anonymous):

oh my goodness..thank you so much!